La présentation est en train de télécharger. S'il vous plaît, attendez

La présentation est en train de télécharger. S'il vous plaît, attendez

SPS LS1 Planning meeting (No.26) 11/12/13 David Mcfarlane

Présentations similaires


Présentation au sujet: "SPS LS1 Planning meeting (No.26) 11/12/13 David Mcfarlane"— Transcription de la présentation:

1 SPS LS1 Planning meeting (No.26) 11/12/13 David Mcfarlane

2 Topics Lifts and Monte-charges. Upcoming works. Maintenance 18 kV. CSA presentation Planning. False floors open in the Bas. E.C.Rs. Christmas Vacation Access rights and safety equipment. Important Information. A.O.B

3 Lift Maintenance Dec/Jan

4 Upcoming works Replacement of Irradiated cables (BA1) – Now until March 2014 Civil Engineering works (TT10) – Now until End of year Replacement of Septas (BA2) – 28 th Oct until end of year Kicker conditioning (BA6) – Dec to mid Jan Re-installation of beam line and equipment in TT10 – January 2014

5 Maintenance 18 kV EMD2*9 JURA du vendredi 20 dec au samedi 21 dec La maintenance 18 kV de la station majeure de JURA ME9 est impérative et critique La maintenance des barres des tranches 1 et 3 en cours actuellement. Elle naffecte pas le fonctionnement de lauto transfert et du réseau secours. Au contraire de la maintenance des barres de la tranche 2 qui nécessitera la mise hors service du système auto transfert et de la liaison Normal/secours. Labsence de la fonction secours ME9 a pour conséquence linterdiction daccès en souterrain LHC1 –ATLAS –LHC18 – SPS6 –SPS7 pendant les 3 à 4 jours de maintenances Par ailleurs des interventions du contractant Schneider E078 sont déjà planifiées W50 et W51 au LHC6 et LHC7… Suite à nos divers échanges de ce jour et au TIOC, nous pouvons proposer le planning suivant Lundi 16 déc. au mercredi 18 déc. : maintenance 18kv LHC7 selon planning LHC Jeudi 19 déc. : jour de repos des techniciens Schneider (et OP) Vendredi 20 déc. au dimanche 22 déc. : maintenance des barres de la tranche 2 JURA : -> pas daccès en souterrain LHC1 –ATLAS –LHC18 – SPS6 –SPS7 A noter quil ny aura pas de coupure utilisateur, les réseaux aval étant réalimentés sans coupure ACCESS INTO BA6 &BA7 WILL BE BLOCKED FROM THURSDAY LUNCHTIME (19 TH DECEMBER)

6 CSA presentation

7 NOTRE MISSION Interface avec les responsables des accès des zones (GS-ASE) Sassurer que les accès aux accélérateurs sont fermés, Faciliter laccès des usagers/intervenants, Sassurer que les intervenants respectent les réglements – Droits daccès, – Port des EPI, – Bonne utilisation PAD/MAD intervenir sur les alarmes Faire remonter les anomalies liées aux accès – Vers TI (problème technique) – Vers GS-DI (non-conformité des personnels / non respect des procédures)

8 NOTRE BESOIN Échanciers des interventions par zone Appeler lagent «ZORA» (n° ) – En début et fin dintervention – Quand une porte est «forcée» dans les accélérateurs – Refermer toutes les portes que vous ouvrez – Autres problèmes lié aux accès

9 EN UndergroundSur BA1 UndergroundSur BA2 UndergroundSur BA3 UndergroundSur BA4 UndergroundSur BA5 UndergroundSur BA6 TI8TI2 TRANSFER TUNNELS UndergroundSur BA7 UndergroundSur BA80/TCC Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. General maintenance Cold check out & Start-up DSO tests/HV tests EPC Commissioning tests Vacuum Ions to NA61 General Maintenance General Maintenance General Maintenance General maintenance General Maintenance UA9 works RF upgrade works EPC works Civil Engineering RF upgrade works RF Upgrade works Survey Horizontal Vertical smoothing Cables Irradiated Cables Cables Re-route Cabling Cables EL Studies Prep work in LSS1 Re-install equipment removed for cable works Survey SPS LS1 schedule Beam on Cool down/RP survey and Magnet tests Protons to the North Area Protons to LHC Cool down NO ACCESS Cool down Limited ACCESS Magnets Beam off : 16/02/13 Beam on: 13/10/14 Open Access : 27/03/13 Close Access : 27/06/14 Fibres CV works CV studies CV Controls upgrade Vacuum TT10 Consolidation 18 KV Septa works Kicker Septa works Kicker TT20 pipework Magnets 18 KV Civil Engineering Magnets Ring-main Vacuum Crab Cavities EPC works Cooling tower 863 Ring-main Cables Fibres 18 KV Civil Engineering in TT10 Fibres CV maintenance Cables Kicker Cables RAL Septa replacement

10 False floors lifted in Surface BAs BAsCabling pulling startCable pulling end BA1 (TS1- works) BA1 (Irradiated cable works) BA1 (Fibre Optics) 29/04/ /10/ /04/ /08/ /02/ /04/2013 BA206/01/201411/03/2014 BA3 BA3 (Fibre Optics) 27/03/ /04/ /05/ /05/2014 BA4 BA4 (Fibre Optics) 27/05/ /10/ /07/ /10/2013 BA5 BA5 (Fibre Optics) 27/03/ /09/ /05/ /10/2013 BA6 BA6 (Fibre Optics) 13/09/ /5/ /11/ /05/2013 BA714/10/201307/11/2013 Below is a table showing the dates that EL will be pulling cables at each point of the SPS. During these periods large sections of the false floor in these surface buildings will be lifted. Any works in the BAs that involve lifting of the false floors MUST be reported to the Coordination team If you require access to racks or other equipment in the surface buildings during these periods you must check with planning first.

11 Engineering Change Requests Templates and Guidelines for Engineering Change Requests: – EDMS: SPS ECR template (in DFS): – https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/A pplications/CERN/Workgroup%20Templates/SPS/ECR-Template- SPS.dotx https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/A pplications/CERN/Workgroup%20Templates/SPS/ECR-Template- SPS.dotx

12 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 1.Removal of AEWB Pick-up from position to (EDMS No ) RELEASED 2.Removal of a BPWA Pick-up from position to (EDMS No ) RELEASED 3.Installation of a new Wideband Kickers system in position (EDMS No ) RELEASED 4.Installation of a new HBW Damper Pick-up in place of BPWA & AEWB (EDMS No ) RELEASED 5.Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2 (EDMS No ) RELEASED 6.Reinstallation of pump and gauges (Pirani + Penning) in (vacuum sector 150) (EDMS No ) RELEASED 7.Modification LSS1 sectorization (new VVSF + displacement VVSB 11772) (EDMS No v.2) RELEASED 8.Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS. (EDMS No ) UNDER IMPLEMENTATION

13 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 9.Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5. (EDMS No ) RELEASED 10.SPS Beam Position Monitoring Prototyping Setup in LSS4. (EDMS No ) RELEASED 11.Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper. (EDMS No ) RELEASED 12.Installation d'un nouveau moniteur de position de type BPCR en LSS2 au (EDMS No ) RELEASED 13.Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel. (EDMS No ) UNDER IMPLEMENTATION 14.Interlocking of Injection Tunnel upstream areas with LHC powering. (EDMS No ) RELEASED 15.Move of BPH to BPH (EDMS No ) RELEASED 16.Optical Fibre Racks. (EDMS No ) UNDER IMPLEMENTATION

14 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 17.Removal of old fixed TMP vacuum groups in (EDMS No ) RELEASED 18.BYBPM racks under the SPS dipole magnets as part of the Upgrade of Beam Position Monitors electronics (MOPOS). (EDMS No ) UNDER IMPLEMENTATION 19.Installation de Coffrets d'Alimentation en LSS1 et TT10. (EDMS No ) RELEASED 20.Installation d'un rack de jonction RX11695=TS1 au BA1, afin de limiter les longueurs des câbles à changer lors d'un futur remplacement. (EDMS No ) UNDER IMPLEMENTATION 21.Installation de nouveaux chemins de câbles en LSS1+ du SPS. (EDMS No v.2) RELEASED 22.TPSN replacement in LSS2 of SPS. (EDMS No v.1) RELEASED 23.Installation d'un nouveau chemin de câbles en LSS2- du SPS. (EDMS No v. 0.1) IN WORK 24.Replacement of pumps in 3 tanks MKP in LSS1 of the SPS. (EDMS No v. 0.1) UNDER APPROVAL

15 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 25.Suppression of VVFA in LSS1 of SPS. (EDMS No v. 0.1) UNDER APPROVAL 26.New shielding wall behind TIDV in SPS. (EDMS No v. 0.1) IN WORK

16 The Christmas Vacation There is no planned works or accesses into the SPS during the 2 week vacation. There are no requests for services during this period?? BA6 and BA7 will be closed from 12:00 on Thursday the 19 th December

17 Access rights and safety equipment Everyone who enters the SPS MUST have the following: – Completed the correct courses. Safety course levels 1,2 and 3. Habilitation Electrique. Radiological Protection. Self Rescue Mask Training. – The correct access authorisation (via EDH). – A valid approved IMPACT. (with their name included on this IMPACT) – They must know the number of this IMPACT. – Safety shoes. – Biocell. – Electronic dosimetre. – Passive dosimetre. – Helmet plus Light. Please note that personnel are not allowed to work in the tunnel alone!!

18 Important information Useful Contact Numbers – SPS Coordination Team David MCFARLANE Franck BAIS – James RIDEWOOD – Control Room SPS CCC – TI CCC – CSA – RP Christophe Tromel – Angelito Herve – Nadine Conan (BA1) – Elodie Aubert Kurt Weiss (BA1) – Coordination Safety Emmanuel Paulat – John Etheridge

19 Please make sure you fill out you IMPACTS correctly!!!! There are currently 19 IMPACTs that are LATE. If you have completed your work then you need to close the related IMPACT. If you have ANY concerns about not being complete before the 27 th June I need to know about it NOW!!!! The SPS cooling towers will be stopped 03/02/14 to 07/02/14 for EL works on switchboard, transformer and SPS pumping station LS2!!!!!! Work has already begun on the planning for LS2. If you are planning on doing any major works in LS2 for consolidation then we need to know about it now. – Please can you send me an stating the main tasks you are planning/hoping to do during LS2 A.O.B.

20 SPS coordination website \\cernhomeJ.cern.ch\J\jocksoft\Public\LS1 Next meeting 08/01/ R-028


Télécharger ppt "SPS LS1 Planning meeting (No.26) 11/12/13 David Mcfarlane"

Présentations similaires


Annonces Google